OpenStudy (anonymous):

Use the distributive law to factor each of the following. Check by multiplying. 13 + 13x

OpenStudy (anonymous):

See, the distributive Law says that: a(b + c) = ab + ac Or, ab + ac = a(b + c) So, you can use this law in the problem and see what you get after applying that???

OpenStudy (anonymous):

im more confused now then i was

OpenStudy (anonymous):

Ha ha ha.. Cool buddy.. Tell me from (13 + 13x), what you can take as common???

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Yes, then do that dear... What you will get????

OpenStudy (anonymous):

Why are you directly giving answers and solutions @Shane_B ???

OpenStudy (shane_b):

I normally don't...but this one was pretty simple. My apologies.

OpenStudy (anonymous):

It is ok dear...

OpenStudy (shane_b):

It's gone :)